# Agentic Program Verification

### v1

@@ 4. Proof Automation with AutoRocq
Overview. Figure 2 shows the overall workflow of AutoRocq. At a high level, it is an LLM agent that takes autonomous actions to carry out a human-like proving process. Specifically, given a lemma and the initial context (i.e., proof state), AutoRocq performs an agentic context analysis to generate a tactic if the context is sufficient, or a query command if additional context is needed (Section 4.1). Then, AutoRocq generates a sequence of tactics by interpreting formal proof representations (i.e., proof tree), analyzing them, and communicating with the Rocq proof assistant autonomously (Section 4.2). During communication, a context-assisted feedback loop is established to either refine an incorrect tactic application or provide additional context to the LLM if the errors persist (Section 4.3). The output of a successful proof is a certificate consisting of a sequence of tactics that can be replayed in the interactive theorem prover to verify the proof.
@@ 4.1. Context-Aware Tactic Generation
Previous works (lewis2020retrieval; rango-icse25; lu2025palm; zhang2024autocoderover) have demonstrated that additional information helps LLMs generate valid tactics. Such additional information, which we refer to as the context, typically include lemmas, definitions, and existing proof steps that are relevant in constructing the proof. Existing approaches, however, tend to augment LLMs with blindly selected context – typically a long list of premises (lu2025palm) or existing proofs (rango-icse25) ordered by a predefined similarity metric – without a good understanding of the current proof state or the progress of the proof. We argue that, before suggesting a new tactic, an intelligent proof system should be able to analyze the current proof state and the progress of the proof to decide if more context is needed. This system should carefully determine what context information is needed and when to add it, so that unnecessary noise is reduced to a minimum. To support this autonomy, an agentic context-aware search mechanism is needed. Such a context-aware approach can significantly reduce the amount of context added to the LLM, which could help the LLM focus on the current proof state and generate a more accurate tactic.
@@
The supported context query commands are listed in Table 1. By default, AutoRocq supports the standard query commands available in the Rocq GUI (CoqIDE), including Search, Print, Locate, About, and Check. These commands enable AutoRocq to retrieve critical context information on demand, significantly aiding the proof process. Among these, the Search command is the most frequently used (see Section 6.2). Its versatility is the key: it allows for discovery based on name patterns, type signatures, and existing lemma names (search-cmd). This makes Search the primary tool for pinpointing relevant declarations without knowing their exact names or forms.
Table 1. List of Supported Context Query Commands in AutoRocq
@@ 4.2. Proof Tree-Aware Interpretation
The key component empowering AutoRocq’s agency is an expressive proof tree representation that enables high-level interpretation of the proof derivation. Compared to the textual or linear representation of proof states (e.g., a list of goals), the explicit tree structure conduces understanding of the proving process. Formally, we define a proof tree as follows (lample2022hypertree; prooftree).
@@
Let $\Sigma$ be the proving context [^1] Here, $\Sigma$ captures both the global environment and the local context, as they are not differentiated in our context search., $G$ be the initial proof goal, and $S$ be a list of tactics that proves $G$ under $\Sigma$. A proof tree $T(\Sigma,G,S)$ is inductively defined as follows:
@@
Intuitively, a proof script $S$ induces a proof tree $T$ in its enclosing context $\Sigma$, rooted at the original proof goal $G$. Each node represents a subgoal to be proved, produced a tactic application is represented as an edge. Leaves are subgoals that are trivially true or could be easily proved with a single tactic. The goal is complete when all the sub-goals in the leaves are proved.
AutoRocq explicitly maintains a (partial) proof tree as it progresses in tactic generation. After a successful tactic application, it examines the open subgoals from the Rocq proof assistant. If the tactic successfully proves the current subgoal (i.e. a leaf node), AutoRocq marks it as closed, and shifts its focus to another residual subgoal. When the tactic creates multiple subgoals, it creates these nodes in the tree and determines one of the new subgoals to work on. In this way, the proof tree is updated by extending the current node. On the other hand, if the tactic fails, our agent remains at the current node on the proof tree for further attempts. During the entire proving process, the proof states in subgoals and tactic applications are continuously updated until a complete tree is constructed. Figure 1(b) shows an example of a proof tree after all tactics were applied in the motivating example presented in Figure 1(a). From the tree, we could see that four subgoals were generated in total, and the lemmas with a violet background were retrieved from our agentic context search. The proof is completed when all leaves are closed.
Whenever the proof tree is updated, AutoRocq interprets the tree structure comprehensively to reevaluate the current progress. It does so by traversing the (partial) proof tree to identify the open subgoals, and how they relate to the existing proof structure. With a holistic view of the entire derivation process, it then proceeds to generate a sequence of tactics iteratively.
We note that our key contribution is not about formulating the proof tree structure. Instead, our insight lies in the realization that high-level interpretation enables LLM agents to carry out autonomous decision-making and actions. The benefits provided by our proof tree representation are twofold. First, it captures the hierarchical structure of the proof, including the relationships between tactics and subgoals, which provides a richer context for interpreting the proof process. Second, it allows AutoRocq to systematically break down the proof into smaller subgoals and tackle them one by one, which is more aligned with how humans approach proving.
@@ 4.3. Context-Assisted Feedback Handling
To emulate a trial-and-error feedback loop, we augment AutoRocq with feedback handling mechanisms to collaborate with the Rocq ITP. Based on the error that occurs, AutoRocq employs two strategies during tactic generation: (1) revising the tactic based on the error message when a single error occurs, and (2) initiating a context search to retrieve additional context when persistent errors occur. In addition, AutoRocq takes positive feedback from past successes. We delegate each mechanism to a subsection as below.
@@ 4.3.2. Handling Persistent Errors with Context Search
4.3.3. Managing Historical Proofs
When a lemma is successfully proved, AutoRocq stores comprehensive tactic history records that capture the complete context and evolution of each proof step. Each record contains the applied tactic, the proof goals before and after tactic application, the available hypotheses and their changes, along with metadata including the theorem name, and the tactic ID within the proof sequence. This rich historical information serves multiple purposes: (1) it enables AutoRocq to learn from successful proof patterns and reuse effective tactic sequences in similar contexts, (2) it provides detailed examples for context-aware tactic generation by showing how specific goals were transformed, and (3) it supports the feedback handling mechanism by offering concrete instances of successful problem-solving strategies. By maintaining this detailed provenance of proof construction, AutoRocq can build a knowledge base of proven tactics that enhances its capability to tackle new lemmas with similar structural patterns or mathematical properties (the experiment results presented in Section 6.2 also support our claim).
Implementation. We implemented AutoRocq in approximately 8k lines of Python code with a modular architecture that separates the core proving logic, theorem prover interface, and supporting utilities. The main Rocq backend is implemented using the CoqPyt (v1.0.0) library (carrott2024coqpyt) to interact with the Rocq proof assistant. The implementation follows an iterative workflow as shown in Figure 2. The modular design enables easy extension and maintenance while supporting the three key components of our approach: context-aware tactic generation, proof tree management, and feedback handling. Historical proof data is stored in JSON format to enable learning from successful proof patterns for future lemmas. For the LLM backend, we use GPT-4.1 as the backbone model for all LLM interactions. We set the temperature to 0 for reproducibility.
@@ 5. Experimental Setup
Is AutoRocq more effective at proving lemmas than state-of-the-art approaches?
@@
5.1. Preparation of Lemmas from Programs
@@
While benchmarks like CoqGym (yang2019coqgym) provide a valuable corpus of human-written theorems, they are drawn primarily from mathematical libraries and do not capture the complexity of program verification lemmas found in real-world software. Since AutoRocq is designed to automate program verification, its evaluation requires a benchmark of lemmas extracted from real programs. To our knowledge, no such dataset currently exists, leaving a critical gap in the program verification field.
To address this gap, we construct a new benchmark by systematically extracting lemmas from real-world C programs. This benchmark enables the evaluation of AutoRocq and facilitates future research in automatic program verification. We source our subject programs from SV-COMP (svcomp) for two key reasons: their widespread adoption in the formal verification community ensures they represent a diverse set of C language constructs, and their inclusion of ground-truth specifications provides meaningful proof obligations. Our selection focused on the most common property types in SV-COMP: functional correctness, defined by the unreachability of error calls, and non-overflow. The final criteria mandated that a program must (1) be verified for both properties and (2) exhibit deterministic behavior (e.g., no multi-threading). Applying these criteria yielded a final benchmark of 131 C programs from SV-COMP. The programs have an average of 43.06 lines of code, with the largest (the BusyBox (busybox) utility basename) comprising 428 lines.
@@ 5.1.2. Lemma Extraction Methodology
We designed a systematic method to automatically extract lemmas from the selected SV-COMP programs. We focus on generating non-trivial lemmas that necessitate reasoning about program semantics, as opposed to simple lemmas that can be discharged with basic tactics (e.g., auto.). Our approach utilizes Frama-C (cuoq2012framac) to generate the necessary proof obligations from the program code.
Technically, we first use the RTE plug-in (rte) to annotate the source code with formal contracts in the ANSI/ISO C Specification Language (ACSL), including preconditions and postconditions. As Frama-C cannot automatically infer loop invariants, we address this by employing LLMs to generate candidate invariants based on the loop’s context and structure. Each candidate is validated through property testing with the Eva plug-in (eva); unsuccessful candidates are refined iteratively until a verifiably correct invariant is established. Note that while property testing can falsify incorrect invariants by discovering counterexamples, it cannot formally prove the correctness of an invariant – that it holds for all possible program executions. We justify this design choice as loop invariant inference is a long-standing, challenging problem (specgen-icse25; learning-invariants) that is orthogonal to our core focus on proof automation. Once the code is fully annotated and verified, we employ Frama-C’s WP plug-in (wp) to generate proof obligations. These obligations are automatically translated into lemmas, which can be discharged in the Rocq proof assistant. The overall framework completes an end-to-end automated workflow from C programs to verifiable Rocq lemmas.
As a result, we collected 641 lemmas extracted from the verification of two types of target properties: non-overflow and functional correctness. Since both property types often require reasoning about loops, a significant portion of the generated proof obligations are related to loop invariants. These are essential intermediate lemmas needed to conclude that either a non-overflow or a functional correctness property holds for the entire program. Consequently, the 641 lemmas comprise three categories, reflecting their role in the proof: i. non-overflow lemmas (203, 31.7%): direct proof obligations for non-overflow properties; ii. functional correctness lemmas (153, 23.9%): direct proof obligations for functional properties; and iii. loop invariant lemmas (285, 44.5%): Supporting lemmas required to prove the loop invariants necessary for establishing both non-overflow and functional correctness properties. Together, these lemmas represent a diverse set of challenging proof obligations that arise in real-world program verification tasks.
@@ 5.1.3. Comparison of Extracted Lemmas and Existing Lemmas
To quantify the complexity and difficulty of lemmas extracted from SV-COMP programs, we compare them against existing lemmas from CoqGym. To do this, one commonly used proxy is the difficulty of proving a lemma, since more sophisticated lemmas generally necessitate longer proofs involving more cases and derivation steps. Unfortunately, this is infeasible for our purpose, since no ground-truth proofs are readily available for our extracted lemmas. As such, we directly examine these lemmas themselves from the extracted 641 lemmas from Section 5.1 and 625 lemmas across four projects (i.e., dblib, zfc, hoare-tut, and huffman) from CoqGym’s testing partition (yang2019coqgym). We choose these four projects as they represent the common type of existing human-written lemmas, such as logical tautology, coding algorithms, and theory formalizations, and they have been extensively studied in the research community (blaauwbroek2020tactician; lu2025palm; rango-icse25). Specifically, we propose two metrics: (a) term count: the number of terms (variables, quantifiers, operators) to gauge structural complexity, and (b) hypothesis count: the number of assumptions to measure the richness of the semantic context.
(a) Histogram of different complexity metrics: lemmas from SV-COMP programs (blue) vs. CoqGym (orange). Not a single lemma from CoqGym involves more than 100 terms or 7 hypotheses.
In 4(a), we report the complexity distribution of the lemmas by measuring the percentage of lemmas (y-axis) with a given complexity metric (term or hypothesis count, x-axis). For both metrics, the lemmas from CoqGym (orange) densely concentrate on the left end of the axis, suggesting uniformly lower complexity. In contrast, lemmas extracted from SV-COMP programs (blue) exhibit a more even distribution, with a lower peak and a long right tail, reflecting generally higher complexity. Concretely, 48% of these lemmas from SV-COMP programs consist of ¿100 terms, and 52% of them are associated with ¿7 hypotheses; no lemma from CoqGym satisfies either constraint. The stark contrast across both metrics suggests that the obligations extracted from real-world programs, which often capture intricate logic in the code, tend to have much higher complexity than their benchmark counterparts. Concretely, below shows an example lemma from hoare_tut, a project in CoqGym. It states that the non-equality function Zneq_bool correctly returns false only if x = y. It is much simpler than lemmas from program verification tasks, as shown in Figure 1(a).
@@
5.2. Baselines and Benchmarks
Baselines. We compare AutoRocq with state-of-the-art tools developed for proof automation in Rocq. We choose four baselines, namely Rango (rango-icse25), PALM (lu2024proof), QEDCartographer (sanchez2024qedcartographer) (or QEDC for short), and Proverbot9001 (sanchez2020generating) (or P9001 for short). These tools have shown exceptional results on proof automation tasks, and they employ different machine-learning techniques: standard LLMs, fine-tuned LLMs, reinforcement learning, and recurrent neural networks (RNNs), respectively. AutoRocq and PALM, which invoke closed-source LLM through APIs, use GPT-4.1 as the backend with temperature 0. For the other three baselines, we use the pre-trained weights made available in their artifacts. These include a fine-tuned version of DeepSeek-Coder used by Rango, a reinforcement-learning-based tactic selection agent employed by QEDC, and a custom tactic prediction network in P9001. They also have access to a single Nvidia A40 GPU with 48GB of VRAM if needed. We use the default settings for all the tools. Rango times out after 10 minutes, and QEDC is limited to 512 generation steps. PALM and P9001 are executed until completion. We set up all the baselines in Docker on Ubuntu 22.04. Other than PALM, which uses older Rocq 8.12 for compatibility reasons, all the other tools are configured with Rocq 8.18.
Benchmarks used in Evaluation. To conduct a comprehensive evaluation, we use the two benchmark sets mentioned in Section 5.1.3, including 625 lemmas from CoqGym (yang2019coqgym) and 641 lemmas extracted from real-world programs in SV-COMP (svcomp). At the end, we have 1,266 lemmas evenly distributed across CoqGym and SV-COMP programs for our evaluation.

### v3

@@ 4. Proof Automation with AutoRocq
Overview. Figure 2 shows the overall workflow of AutoRocq. At a high level, it is an LLM agent that takes autonomous actions to carry out a human-like proving process. Specifically, given a lemma and the initial context (i.e., proof state), AutoRocq performs an agentic context analysis to generate a tactic if the context is sufficient, or a query command if additional context is needed (Section 4.1). Then, AutoRocq generates a sequence of tactics by interpreting formal proof representations (i.e., proof tree), analyzing them, and communicating with the Rocq proof assistant autonomously (Section 4.2). During communication, a context-assisted feedback loop is established to either refine an incorrect tactic application or provide additional context to the LLM if the errors persist, and a on-the-fly learning component is proposed to use successful proofs to help prove new lemmas (Section 4.3). The output of a successful proof is a certificate consisting of a sequence of tactics that can be replayed in the interactive theorem prover to verify the proof.
@@ 4.1. Context-Aware Tactic Generation
Previous works (Lewis et al., 2020; Thompson et al., 2025; Lu et al., 2024; Zhang et al., 2024) have demonstrated that additional information helps LLMs generate valid tactics. Such additional information, which we refer to as the context, includes lemmas, definitions, and existing proof steps that are relevant in constructing the proof. Existing approaches, however, tend to augment LLMs with blindly selected context – typically a long list of premises (Lu et al., 2024) or existing proofs (Thompson et al., 2025) ordered by a predefined similarity metric – without a good understanding of the current proof state or the progress of the proof. We argue that, before suggesting a new tactic, an intelligent proof system should be able to analyze the current proof state and the progress of the proof to decide if more context is needed. This system should carefully determine what context information is needed and when to add it, so that unnecessary noise is reduced to a minimum. To support this autonomy, an agentic context-aware search mechanism is needed. Such a context-aware approach can significantly reduce the amount of context added to the LLM, which could help the LLM focus on the current proof state and generate a more accurate tactic.
@@
The supported context query commands are listed in Table 1 and the examples of query output are shown in Table 2. By default, AutoRocq supports the standard query commands available in the Rocq GUI (CoqIDE), including Search, Print, Locate, About, and Check. These commands enable AutoRocq to retrieve critical context information on demand, significantly aiding the proof process. Among these, the Search command is the most frequently used (see Section 6.3). Its versatility is the key: it allows for discovery based on name patterns, type signatures, and existing lemma names (Rocq Development Team, 2025). This makes Search the primary tool for pinpointing relevant declarations without knowing their exact names or forms.
Table 1. List of supported context query commands in AutoRocq.
@@
Table 2. Examples of context query commands and their outputs.
Example Command Output (reduced version)
Search (Z.quot _ _). Zquot.Z_mult_quot_le: forall a : int, 0 ÷ a = 0
Z.quot_unique_exact: forall a : int, a ÷ 0 = 0
Print is_sint32. is_sint32=fun x:int => -2147483648 <= x < 2147483648...
Locate to_sint64. Constant main_assert.to_sint64.
About Z.abs. Z.abs : int -> int.
Check why_decidable_eq. why_decidable_eq : forall x y : bool, x = y + x <> y
@@ 4.2. Proof Tree-Aware Interpretation
The key component empowering AutoRocq’s agency is an expressive proof tree representation that enables high-level interpretation of the proof derivation. Compared to the textual or linear representation of proof states (e.g., a list of goals), the explicit tree structure conduces understanding of the proving process. Formally, we define a proof tree as follows (Lample et al., 2022; Hendrik, 2025).
@@
Let $\Sigma$ be the proving context [^2] Here, $\Sigma$ captures both the global environment and the local context, as they are not differentiated in our context search., $G$ be the initial proof goal, and $S$ be a sequence of tactics that transforms $G$ under $\Sigma$. A proof tree $T(\Sigma,G,S)$ is inductively defined as:
@@
Note that the $S$ may or may not be a complete tactic sequence that proves $G$. Intuitively, a proof script $S$ induces a proof tree $T$ in its enclosing context $\Sigma$, rooted at the original proof goal $G$. Each node represents a subgoal to be proved, and a tactic application is represented as an edge. Leaves are subgoals that are trivially true or could be easily proved with a single tactic. The goal is complete when all the sub-goals in the leaves are proved.
AutoRocq explicitly maintains a proof tree as it progresses in tactic generation. After a successful tactic application, it examines the open subgoals from the Rocq proof assistant. If the tactic successfully proves the current subgoal (i.e., a leaf node), AutoRocq marks it as closed, and shifts its focus to another residual subgoal. When the tactic creates multiple subgoals, it creates these nodes in the tree and determines one of the new subgoals to work on. In this way, the proof tree is updated by extending the current node. On the other hand, if the tactic fails, our agent remains at the current node on the proof tree for further attempts. During the entire proving process, the proof states in subgoals and tactic applications are continuously updated until a complete tree is constructed. Figure 1(b) shows an example of a proof tree after all tactics were applied in the motivating example presented in Figure 1(a). From the tree, we could see that four subgoals were generated in total, and the lemmas with a violet background were retrieved from our agentic context search. The proof is completed when all leaves are closed.
Whenever the proof tree is updated, AutoRocq interprets the tree structure comprehensively to reevaluate the current progress. It does so by traversing the proof tree to identify the open subgoals and how they relate to the existing proof structure. With a holistic view of the entire derivation process, it then proceeds to generate a sequence of tactics iteratively.
We note that our key contribution is not about formulating the proof tree structure. Instead, our insight lies in the realization that high-level interpretation enables LLM agents to carry out autonomous decision-making and actions. The benefits provided by our proof tree representation are twofold. First, it captures the hierarchical structure of the proof, including the relationships between tactics and subgoals, which provides a richer context for interpreting the proof process. Second, it allows AutoRocq to systematically break down the proof into smaller subgoals and tackle them one by one, which is aligned with how humans approach proving (Scherer, 2017; Shi et al., 2025).
@@ 4.3. Context-Assisted Feedback Handling
To emulate a trial-and-error feedback loop, we augment AutoRocq with feedback handling mechanisms to collaborate with the Rocq ITP. Based on the error that occurs, AutoRocq employs two strategies during tactic generation: (1) revising the tactic based on the error message when a single error occurs, and (2) initiating a context search to retrieve additional context when persistent errors occur. In addition, AutoRocq takes positive feedback from past successes. We delegate each mechanism to a subsection as follows.
@@ 4.3.2. Handling Persistent Errors with Context Search
4.3.3. Managing Historical Proofs for On-the-fly Learning
AutoRocq learns on the fly by maintaining an archive of successful proofs. This archive is accumulated throughout a proving campaign, and AutoRocq adapts its strategy when proving new lemmas. Specifically, when a lemma is successfully proved, AutoRocq stores comprehensive tactic history records that capture the complete context and evolution of each proof step. Each record contains the applied tactic, the proof goals before and after tactic application, the available hypotheses and their changes, along with metadata including the theorem name, and the tactic ID within the proof sequence. This rich historical information serves multiple purposes: (1) it enables AutoRocq to learn from successful proof patterns and reuse effective tactic sequences in similar contexts, (2) it provides detailed examples for context-aware tactic generation by showing how specific goals were transformed, and (3) it supports the feedback handling mechanism by offering concrete instances of successful problem-solving strategies. By maintaining this detailed provenance of proof construction, AutoRocq can build a knowledge base of proven tactics that enhances its capability to tackle new lemmas with similar structural patterns or mathematical properties (the experiment results presented in Section 6.3 also support our claim).
4.4. Implementation
We implemented AutoRocq in approximately 8k lines of Python code with a modular architecture that separates the core proving logic, theorem prover interface, and supporting utilities. The main Rocq backend is implemented using the CoqPyt (v1.0.0) library (Carrott et al., 2024) to interact with the Rocq proof assistant. The implementation follows an iterative workflow as shown in Figure 2. The modular design enables easy extension and maintenance while supporting the three key components of our approach: context-aware tactic generation, proof tree management, and feedback handling. Historical proof data is stored in JSON format to enable learning from successful proof patterns for future lemmas. For the LLM backend, we use GPT-4.1 as the backbone model for all LLM interactions. We set the temperature to 0 for reproducibility.
@@ 5. Experimental Setup
Is AutoRocq effective at proving mathematical lemmas from CoqGym?
@@
Is AutoRocq effective at proving lemmas from program verification tasks?
RQ.3::
@@
5.1. Preparation of Benchmark Lemmas
5.1.1. Lemmas from Existing Datasets
CoqGym (Yang and Deng, 2019) and CoqStoq (Thompson et al., 2025) are two well-known datasets for evaluating neural theorem provers in Rocq, and have been extensively studied in the community (Blaauwbroek et al., 2020; Lu et al., 2024; Thompson et al., 2025). For fair and comprehensive evaluation, we include only the lemmas that come from the intersection of CoqStoq and CoqGym testing sets. This is because other projects from CoqGym’s test set are included in CoqStoq’s training set, and thus may cause inflated results. Specifically, these include seven projects (i.e., dblib, zfc, hoare-tut, huffman, buchberger, PolTac, and zorns-lemma), which represent common mathematical and human-written lemmas, such as logical tautology, coding algorithms, and theory formalizations. In total, there are 1,717 mathematical lemmas selected to evaluate comparative approaches.
@@ 5.1.1. Lemmas from Existing Datasets
While CoqGym (Yang and Deng, 2019) and CoqStoq (Thompson et al., 2025) provide a valuable corpus of human-written theorems, they are drawn primarily from mathematical libraries and do not capture the complexity of program verification lemmas found in real-world software. Since AutoRocq is designed to automate program verification, its evaluation requires a benchmark of lemmas extracted from real programs. To our knowledge, no such dataset currently exists, leaving a critical gap in the program verification field.
To address this gap, we construct a new benchmark by systematically extracting lemmas from real-world C programs. This benchmark enables the evaluation of AutoRocq and facilitates future research in automatic program verification. We source our subject programs from SV-COMP (Beyer and Strejček, 2025) for two key reasons: their widespread adoption in the formal verification community ensures they represent a diverse set of C language constructs, and their inclusion of ground-truth specifications provides meaningful proof obligations. Our selection focused on the most common property types in SV-COMP: functional correctness, defined by the unreachability of error calls, and non-overflow. The final criteria mandated that a program must (1) be verified for both properties and (2) exhibit deterministic behavior (e.g., no multi-threading). Applying these criteria yielded a final benchmark of 131 C programs from SV-COMP. The programs have an average of 43.06 lines of code, with the largest (the utility basename in BusyBox (7)) comprising 428 lines.
@@ 5.1.3. Lemma Extraction Methodology
We designed a systematic method to automatically extract lemmas from the selected SV-COMP programs. We focus on generating non-trivial lemmas that necessitate reasoning about program semantics, as opposed to simple lemmas that can be discharged with basic tactics (e.g., auto.). Our approach utilizes Frama-C (Kirchner et al., 2015) to generate the necessary proof obligations from the program code.
Technically, we first use the RTE plug-in (54) to annotate the source code with formal contracts in the ANSI/ISO C Specification Language (ACSL), including preconditions and postconditions. As Frama-C cannot automatically infer loop invariants, we address this by employing LLMs to generate candidate invariants based on the loop’s context and structure. Each candidate is validated through property testing with the Eva plug-in (17); unsuccessful candidates are refined iteratively until a verifiably correct invariant is established. Note that while property testing can falsify incorrect invariants by discovering counterexamples, it cannot formally prove the correctness of an invariant – that it holds for all possible program executions. We justify this design choice as loop invariant inference is a long-standing, challenging problem (Ma et al., 2025; Garg et al., 2016) that is orthogonal to our core focus on proof automation. Once the code is fully annotated and verified, we employ Frama-C’s WP plug-in (Frama-C Developers, 2026) to generate proof obligations. These obligations are automatically translated into lemmas, which can be discharged in the Rocq proof assistant. The overall framework completes an end-to-end automated workflow from C programs to verifiable Rocq lemmas.
As a result, we collected 641 lemmas extracted from the verification of two types of target properties: non-overflow and functional correctness. Since both property types often require reasoning about loops, a significant portion of the generated proof obligations are related to loop invariants. These are essential intermediate lemmas needed to conclude that either a non-overflow or a functional correctness property holds for the entire program. Consequently, the 641 lemmas comprise three categories, reflecting their role in the proof: (1) non-overflow lemmas (203, 31.7%): direct proof obligations for non-overflow properties; (2) functional correctness lemmas (153, 23.9%): direct proof obligations for functional properties; and (3) loop invariant lemmas (285, 44.5%): Supporting lemmas required to prove the loop invariants necessary for establishing both non-overflow and functional correctness properties. Together, these lemmas represent a diverse set of challenging proof obligations that arise in real-world program verification tasks.
@@ 5.1.4. Comparison of Extracted Lemmas and Existing Lemmas
To quantify the complexity and difficulty of verification lemmas extracted in Section 5.1.3, we compare them against existing mathematical lemmas pooled in Section 5.1.1. To do this, one commonly used proxy is the difficulty of proving a lemma, since more sophisticated lemmas generally necessitate longer proofs involving more cases and derivation steps. Unfortunately, this is infeasible for our purpose, since no ground-truth proofs are readily available for our extracted lemmas. As such, we directly examine these lemmas themselves. Specifically, we propose two metrics: (a) term count: the number of terms (variables, quantifiers, operators) to gauge structural complexity, and (b) hypothesis count: the number of assumptions to measure the richness of the semantic context.
(a) Term Count
(b) Hypothesis Count
Figure 3. Histogram of different complexity metrics: lemmas from SV-COMP programs (blue) vs. CoqGym (orange). Less than 1% of lemmas from CoqGym involve more than 100 terms or 7 hypotheses.
In Figure 3, we report the complexity distribution of the lemmas by measuring the percentage of lemmas (y-axis) with a given complexity metric (term or hypothesis count, x-axis). For both metrics, the lemmas from CoqGym (orange) densely concentrate on the left end of the axis, suggesting uniformly lower complexity. In contrast, lemmas extracted from SV-COMP programs (blue) exhibit a more even distribution, with a lower peak and a long right tail, reflecting generally higher complexity. Concretely, 48% of these lemmas from SV-COMP programs consist of ¿100 terms, and 52% of them are associated with ¿7 hypotheses; $1\%$ mathematical lemmas satisfy either constraint. The stark contrast across both metrics suggests that the obligations extracted from real-world programs, which often capture intricate logic in the code, tend to have much higher complexity than their benchmark counterparts. Below shows an example lemma from hoare_tut, a project in CoqGym. It states that the non-equality function Zneq_bool correctly returns false only if x = y. It is much simpler than lemmas from program verification tasks, as shown in Figure 1(a).
@@
In summary, we use the two benchmark sets for our evaluation in Section 6, namely (1) 1,717 mathematical lemmas from the intersection of the test partition of CoqGym (Yang and Deng, 2019) and CoqStoq (Thompson et al., 2025), and (2) 641 verification-derived lemmas from SV-COMP programs (Beyer and Strejček, 2025). In total, we have 2,358 lemmas across different domains.
5.2. Comparative Baselines
We compare AutoRocq with state-of-the-art tools developed for proof automation in Rocq. We choose five baselines, namely Rango (Thompson et al., 2025), PALM (Lu et al., 2024), Copra (Thakur et al., 2024), QEDCartographer (Sanchez-Stern et al., 2025) (or QEDC for short), and Proverbot9001 (Sanchez-Stern et al., 2020) (or P9001 for short). These tools have shown exceptional results on proof automation tasks, and they employ different machine-learning techniques: standard LLMs, fine-tuned LLMs, reinforcement learning, and recurrent neural networks (RNNs), respectively. AutoRocq, PALM, and Copra, which invoke closed-source LLM through APIs, use GPT-4.1 as the backend with temperature 0. For the other three baselines, we use the pre-trained weights made available in their artifacts. These include a fine-tuned version of DeepSeek-Coder used by Rango, a reinforcement-learning-based tactic selection agent employed by QEDC, and a custom tactic prediction network in P9001. They also have access to a single Nvidia A40 GPU with 48GB of VRAM if needed. We use the default settings for all the tools. Rango and Copra times out after 10 minutes, and QEDC is limited to 512 generation steps. PALM and P9001 are executed until completion. We set up all the baselines in Docker on Ubuntu 22.04. Other than PALM, which uses older Rocq 8.12 for compatibility reasons, all the other tools are configured with Rocq 8.18. We delay the discussions on the use of different models and Rocq versions to Section 6.5.